OpenStudy (anonymous):

Simplify if cannot be simplified write on simplest form: 2e + 4f - e + 2f

OpenStudy (anonymous):

e + 6f

OpenStudy (anonymous):

Thx for the medal :D

OpenStudy (anonymous):

Np thxs for helping me

jimthompson5910 (jim_thompson5910):

Combine like terms 2e - e = 2e - 1e = 1e = e 4f + 2f = 6f So that explains how he got e + 6f